# Unconditional correctness of recent quantum algorithms for factoring and computing discrete logarithms

### v1

@@ 2.4. Bounding the contribution of large-order characters
If $t\geqslant 2d$, the bound 12 is trivially satisfied as the right-hand side is $\gg N$. We may thus assume that ${t_{0}\leqslant t\leqslant 2d}$, in which case Lemma 2.12 applies.
Let $E_{t}$ be the set of all characters $\psi$ modulo $N$ such that $L(s,\psi)$ has a zero in the rectangle defined in 10. By Proposition 2.10, this set $E_{t}$ has size
@@
For every $\chi\in\widehat{G}$ of order $\geqslant e^{10d}$, if $\mathbb{E}\mathopen{}\mathclose{{}\left[|F_{\chi,H}(\mathbf{b}_{1},\ldots,\mathbf{b}_{d})|^{2}}\right]>e^{-2td}H^{2d}$ then by Lemma 2.12 we can write
@@
for some integer $0<h<e^{3t}$ and some characters $\psi_{1},\psi_{2}\in E_{t}$. Hence, the left-hand side of 12 is
@@
Since $\widehat{G}^{{M_{*}}}$ and $\widehat{G^{M_{*}}}$ are isomorphic, Lemma 2.5 implies that
(15) |$\mathopen{}\mathclose{{}\left|\big{\{}\chi\in\widehat{G}^{M_{*}}\ :\ \chi^{h}=\psi_{1}\overline{\psi_{2}}\big{\}}}\right|\leqslant h^{d/10}\leqslant e^{3td/10}.$
Inserting 13 and 15 into 14, we conclude that the left-hand side of 12 is
@@
Let $N,G,d,X$ and $\mathbf{b}_{i}$ be as in Lemma 2.11. Let ${M_{*}}\geqslant 1$ be the integer defined in Lemma 2.4. For $H\geqslant e^{10d}$, we have
@@
By Lemma 2.11, we know that $\mathbb{E}\mathopen{}\mathclose{{}\left[|F_{\chi,H}(\underline{\mathbf{b}})|^{2}}\right]^{1/2}\leqslant Cd^{-5d}H^{d}$ for all characters $\chi\in\widehat{G}$ of order $\geqslant e^{10d}$, where $C>0$ is an absolute constant (as before, $\underline{\mathbf{b}}$ stands for $\mathbf{b}_{1},\ldots,\mathbf{b}_{d}$). Therefore,
@@
where $m_{0}:=\lfloor 5d\log d-\log C+1\rfloor$. We may assume that $m_{0}\geqslant\max(t_{0}d,4d\log d)$, since otherwise $d\ll 1$ and Proposition 2.14 is trivially satisfied (given that $|\widehat{G}^{M_{*}}|\leqslant N\ll 1$ when $d\ll 1$).
Applying Lemma 2.13 with $t:=m/d$, we obtain that for every $m\geqslant t_{0}d$,
@@ 2.5. Geometry of numbers
Let $L\geqslant 1$ be an integer. Cover the cube $[-L,L]^{d}$ by $(2L)^{d}$ cubes of side length $1$ in the obvious way. Label these unit cubes $C_{1},\ldots,C_{(2L)^{d}}$ (in any order). Let $V\subset\mathbb{R}^{d}$ be a hyperplane through the origin. Then the number of unit cubes $C_{i}$ intersecting $V$ is
$\leqslant(d+1)(2L)^{d-1}.$
@@
Let $L=H_{1}/H_{0}$. The large cube $[-H_{1},H_{1}]^{d}$ can be covered by $(2L)^{d}$ axis-parallel cubes of side length $H_{0}$ in the natural way. By Lemma 2.15, we can bound
@@
Plugging in our lattice point estimate 16, we get
@@
The linearly independent vectors given by Lemma 2.16 can be upgraded to a genuine basis for $\Lambda$ by standard geometry of numbers, namely Mahler’s theorem. We state this fact in a slightly more general situation.
@@
Let $d\geqslant 1$. Let $\Lambda_{1},\Lambda_{2}\subset\mathbb{R}^{d}$ be full-rank lattices such that $M\Lambda_{1}\subset\Lambda_{2}$ for some integer $M\geqslant 1$. Suppose that $\Lambda_{1}$ contains $d$ linearly independent vectors in $[-H,H]^{d}$ for some $H>0$. Then, $\Lambda_{2}$ admits a basis where each basis vector has Euclidean norm $\leqslant d^{3/2}MH$.
@@
Let $B\subset\mathbb{R}^{d}$ be the unit ball for the Euclidean norm. Let ${0<\lambda_{1}\leqslant\cdots\leqslant\lambda_{d}}$ be the successive minima [^8] See [22, Definition 3.29] for the definition of successive minima. of $B$ with respect to $\Lambda_{2}$. Since $Mv_{1},\ldots,Mv_{d}\in\Lambda_{2}$ are linearly independent, we deduce from the above that $\lambda_{d}\leqslant\sqrt{d}MH$.
By Mahler’s theorem (see [22, Theorem 3.34]), we conclude that $\Lambda_{2}$ admits a basis of vectors of Euclidean norm
$\leqslant d\lambda_{d}\leqslant d^{3/2}MH,$
which is what we needed to show. ∎
@@ 2.6. Short basis vectors
Then, with probability $1-O\big{(}d^{-d}\big{)}$, the lattice
:=$\mathcal{L}:=\Big{\{}(e_{1},\ldots,e_{d},f_{1},\ldots,f_{r})\in\mathbb{Z}^{d+r}\ :\ \prod_{i=1}^{d}\mathbf{b}_{i}^{e_{i}}\prod_{i=1}^{r}\mathbf{x}_{i}^{f_{i}}\equiv 1\!\!\pmod{N}\Big{\}}$
@@
The constant $42$ in the exponent is by no means the limit our techniques. For example, using smooth cutoffs in Definition 2.6 would significantly reduce the error term in Lemma 2.8 and hence lower this constant. We have not performed such optimisations to keep the paper as simple as possible.
Proof of Theorem 2.18.
Let $M:={M_{*}}(N)$ be the integer defined in Lemma 2.4. We introduce the auxiliary lattice
@@
By Lemmas 2.7 and 2.8, we have the lattice point estimate
@@
Applying Proposition 2.14, we deduce that, for all $H\geqslant e^{10d}$,
@@
Let $H_{0}:=d\lceil e^{31(d+r)}\rceil$ and $H_{1}:=\lceil d(d+r)(5/2)^{d+r}\rceil H_{0}$. We apply 17 and 18 twice, once with $H=H_{0}$ and once with $H=H_{1}$ to obtain the following: with probability ${1-O\big{(}d^{-d}\big{)}}$, the two estimates
@@
We now apply Lemma 2.16. By our choice of $H_{0}$ and $H_{1}$, the inequality in the statement is satisfied provided that $d$ is sufficiently large. We can assume that $d$ is large enough as, for $d\ll 1$, we have $N\ll 1$ and Theorem 2.18 is trivially true. Thus, Lemma 2.16 implies that ${\mathcal{L}_{M}\cap[-H_{1},H_{1}]^{d+r}}$ contains $d+r$ linearly independent vectors, with probability ${1-O\big{(}d^{-d}\big{)}}$. By Lemma 2.17, since $M\mathcal{L}\subset\mathcal{L}_{M}$, we conclude that, with probability ${1-O\big{(}d^{-d}\big{)}}$, $\mathcal{L}$ admits a basis of vectors of Euclidean norm
@@
using the bound $M\leqslant e^{10d}$ given by Lemma 2.4. This completes the proof. ∎
@@ 3. Applications to quantum computing
In this section, we prove the correctness of efficient quantum algorithms for factoring and for the discrete logarithm problem by applying our version of Regev’s number-theoretic conjecture, Theorem 2.18.
@@ 3.1. Preparatory lemmas
This is similar to [18, p5] or [17, Lemma 5.6] but for a general value of $m$. Without loss of generality, assume that $d$ is a power of $2$, say $d=2^{l}$. We proceed to compute this product in a binary tree fashion. Let $T(k)$ be the complexity of multiplying any $k$ of these numbers $a_{i}^{t_{i}}$ modulo $N$. Note that $T(2k)\leqslant 2T(k)+O(M(mk))$ where $M(x)$ is the time needed to multiply two integers having at most $x$ bits. By the work of Harvey and van der Hoeven [10], it is known that $M(x)=O(x\log x)$, which leads to the bound
@@
This well-known fact is explained in [17, Section A.1]. ∎
@@
Let $\mathbf{n}_{1},\ldots,\mathbf{n}_{k}$ be i.i.d. random random variables uniformly distributed in $\{1,\ldots,X\}$. Then, the probability that at least $d$ of these $\mathbf{n}_{i}$ are prime numbers not dividing $N$ is $\geqslant 1-O(1/N)$.
@@
$\mathbb{P}\bigg{(}\bigcup_{\begin{subarray}{c}I\subset[k]\\ |I|>k-d\end{subarray}}\bigcap_{i\in I}E_{i}\bigg{)}\leqslant\sum_{l<d}{k\choose l}\mathopen{}\mathclose{{}\left(1-\frac{c}{\log X}}\right)^{k-l}\leqslant dk^{d}e^{-c(k-d)/\log X}\ll e^{-d^{2}}$
@@ 3.2. The discrete logarithm problem
We now prove Theorem 1.2. For convenience, we restate it here.
@@
$\mathcal{L}_{\underline{\mathbf{b}},g,y}:=\Big{\{}(e_{1},\ldots,e_{d},f_{1},f_{2})\in\mathbb{Z}^{d+2}\ :~\ \Bigg{(}\prod_{i=1}^{d}\mathbf{b}_{i}^{e_{i}}\Bigg{)}g^{f_{1}}y^{f_{2}}\equiv 1\!\!\pmod{N}\Big{\}}.$
By Theorem 2.18, with probability $1-O(d^{-d})$, this lattice has a basis consisting of vectors of Euclidean norm $\ll e^{42d}$.
Computing the discrete logarithm of $y$ with respect to the base $g$ reduces (with a polynomial-time classical algorithm) to computing a short basis for $\mathcal{L}_{\underline{\mathbf{b}},g,y}$. To see why this is the case, suppose that we managed to compute a basis $v_{1},\ldots,v_{d+2}$ for $\mathcal{L}_{\underline{\mathbf{b}},g,y}$ with $\max_{i}\mathopen{}\mathclose{{}\left\|v_{i}}\right\|_{2}\ll e^{42d}$. It is then easy to find a vector in $\mathcal{L}_{\underline{\mathbf{b}},g,y}$ of the form $(0,\ldots,0,x,1)$ for some integer $x$ (note that such a vector exists since we assume that $y\in\mathopen{}\mathclose{{}\left\langle g}\right\rangle$). Indeed, this amounts to solving a linear system with integer coefficients. The complexity of solving an integer linear system is polynomial in the dimensions of the matrix and the number of bits of the coefficients, which are both $O(d)$ since $\max_{i}\mathopen{}\mathclose{{}\left\|v_{i}}\right\|_{2}\ll e^{42d}$ (see [3]). This yields an integer $x$ such that $g^{x}\equiv y\pmod{N}$.
@@
    Generate $d$ primes $b_{1},\ldots,b_{d}$ independently and uniformly at random in the set of primes $\leqslant X$ not dividing $N$. To do this, it suffices to generate $d^{4}$ independent random integers $\leqslant X$. By Lemma 3.3, the required conditions will be satisfied for at least $d$ of those with probability $1-O(1/N)$. This first step takes polynomial time on a classical computer using the AKS primality test [1].
@@
    Use the procedure described by Ekerå and Gärtner [7] to obtain a basis for the lattice $\mathcal{L}_{\underline{b},g,y}$. This involves making $O(d)$ calls to a quantum circuit, followed by polynomial-time classical post-processing. This step is now guaranteed to succeed with probability $\Theta(1)$, using the fact that $\mathcal{L}_{\underline{b},g,y}$ has a basis of vectors of Euclidean norm $\ll e^{42d}$ with probability $1-O(d^{-d})$ (Theorem 2.18).
@@
The only modification needed to the analysis of the quantum circuit in [7] comes from the fact that $b_{1},\ldots,b_{d}$ are not quite as small as in [7]. In [7] (and more generally in [17, 18]), the primes $b_{i}$ are assumed to have $O(\log d)$ bits. In the present situation, we instead have the bound $b_{i}\leqslant X=d^{10^{3}d}$, i.e. each $b_{i}$ has $O(d\log d)$ bits.
The only place in [7] where the assumption on the size of the $b_{i}$’s are used is in [7, Lemma 3]. In turn, the only point in the proof of [7, Lemma 3] where this assumption is needed in when [17, Lemma 4.1] is invoked (as a black box). The proof of [17, Lemma 4.1] is given in [17, Section 5], and the size assumption on the $b_{i}$’s only comes up in [17, Lemma 5.6].
The result [17, Lemma 5.6] essentially [^10] There are extra details pertaining to the precise use of qubits (e.g. restoring the ancilla qubits to $|0\rangle$), but these will be the same in our setup. states that there is a quantum circuit using $O(d\log^{3}d)$ gates and $O(d\log^{3}d)$ qubits to perform the computation of $\prod_{i=1}^{d}a_{i}^{t_{i}}$ where $t_{i}\in\{0,1\}$ and $a_{i}$ are integers on $O(\log d)$ bits. This is a special case of Lemma 3.1 (used together with Lemma 3.2) applied with $m=O(\log d)$. In our case, we just apply Lemma 3.1 with $m=O(d\log d)$, together with Lemma 3.2 to convert the classical circuit into a quantum one. This yields a quantum circuit having $O(d^{2}\log^{3}d)$ gates and $O(d^{2}\log^{3}d)$ qubits to compute $\prod_{i=1}^{d}b_{i}^{t_{i}}$.
The number of gates of the quantum circuit [17, Lemma 4.1] is
@@
(because [17, Lemma 5.6] is used $O(d)$ times, see [17, Algorithm 5.2] and the surrounding explanations). Note that this is the same as for Regev’s algorithm (see [18, p5]). In our case, the number of gates is
@@
The space optimisations of Ragavan and Vaikuntanathan keep the total number of qubits for their circuit [17, Lemma 4.1] under
@@
due to the way the qubits are used and restored in the main loop of [17, Algorithm 5.2]. In our situation, the number of qubits is
@@
In summary, our final quantum circuit has $O(n^{3/2}\log^{3}n)$ gates and $O(n\log^{3}n)$ qubits. As noted in [7], the fact that the two elements $g$ and $y$ are not small ($g$ and $y$ can be as large as $N$, as opposed to the $b_{i}$’s) does not affect these complexity bounds. ∎
@@
In the proof of Theorem 1.2, we used the naive Lemma 3.2 to convert a classical circuit into a quantum one. As mentioned in Remark 1.3, it is possible to save a factor of $\log n$ in the space cost for Theorems 1.2 and 1.1 by reusing certain qubits when performing the quantum computation corresponding to Lemma 3.1.

### v2

@@ 3.3. Bounding the contribution of large-order characters
If $t\geqslant 2d$, the bound ˜18 is trivially satisfied as the right-hand side is $\gg N$. We may thus assume that $t_{0}\leqslant t\leqslant 2d$}, in which case Lemma˜3.12 applies.
Let $E_{t}$ be the set of all characters $\psi$ modulo $N$ such that $L(s,\psi)$ has a zero in the rectangle defined in ˜16. By Proposition˜3.10, this set $E_{t}$ has size
@@
For every $\chi\in\widehat{G}$ of order $\geqslant e^{10d}$, if $\mathbb{E}\mathopen{}\mathclose{{\left[|F_{\chi,H}(\mathbf{b}_{1},\ldots,\mathbf{b}_{d})|^{2}}}\right]>e^{-2td}H^{2d}$ then by Lemma˜3.12 we can write
@@
for some integer $0<h<e^{3t}$ and some characters $\psi_{1},\psi_{2}\in E_{t}$. Hence, the left-hand side of ˜18 is
@@
Since $\widehat{G}^{{M_{*}}}$ and $\widehat{G^{M_{*}}}$ are isomorphic, Lemma˜3.5 implies that
(21) $\mathopen{}\mathclose{{\left|\big\{\chi\in\widehat{G}^{M_{*}}\ :\ \chi^{h}=\psi_{1}\overline{\psi_{2}}\big\}}}\right|\leqslant h^{d/10}\leqslant e^{3td/10}.$
Inserting ˜19 and 21 into ˜20, we conclude that the left-hand side of ˜18 is
@@
Let $N,G,d,X$ and $\mathbf{b}_{i}$ be as in Lemma˜3.11. Let ${M_{*}}\geqslant 1$ be the integer defined in Lemma˜3.4. For $H\geqslant e^{10d}$, we have
@@
By Lemma˜3.11, we know that $\mathbb{E}\mathopen{}\mathclose{{\left[|F_{\chi,H}(\underline{\mathbf{b}})|^{2}}}\right]^{1/2}\leqslant Cd^{-5d}H^{d}$ for all characters $\chi\in\widehat{G}$ of order $\geqslant e^{10d}$, where $C>0$ is an absolute constant (as before, $\underline{\mathbf{b}}$ stands for $\mathbf{b}_{1},\ldots,\mathbf{b}_{d}$). Therefore,
@@
where $m_{0}:=\lfloor 5d\log d-\log C+1\rfloor$. We may assume that $m_{0}\geqslant\max(t_{0}d,4d\log d)$, since otherwise $d\ll 1$ and Proposition˜3.14 is trivially satisfied (given that $|\widehat{G}^{M_{*}}|\leqslant N\ll 1$ when $d\ll 1$).
Applying Lemma˜3.13 with $t:=m/d$, we obtain that for every $m\geqslant t_{0}d$,
@@ 3.4. Geometry of numbers
Let $L\geqslant 1$ be an integer. Cover the cube $L,L]^{d}$ by $(2L)^{d}$ cubes of side length $1$ in the obvious way. Label these unit cubes $C_{1},\ldots,C_{(2L)^{d}}$ (in any order). Let $V\subset\mathbb{R}^{d}$ be a hyperplane through the origin. Then the number of unit cubes $C_{i}$ intersecting $V$ is at most $(d+1)(2L)^{d-1}$.
@@
Let $L=H_{1}/H_{0}$. The large cube $H_{1},H_{1}]^{d}$ can be covered by $(2L)^{d}$ axis-parallel cubes of side length $H_{0}$ in the natural way. By Lemma˜3.15, we can bound
@@
Plugging in our lattice point estimate ˜22, we get
@@
The linearly independent vectors given by Lemma˜3.16 can be upgraded to a genuine basis for $\Lambda$ by standard geometry of numbers, namely Mahler’s theorem. We state this fact in a slightly more general situation.
@@
Let $d\geqslant 1$. Let $\Lambda_{1},\Lambda_{2}\subset\mathbb{R}^{d}$ be full-rank lattices such that $M\Lambda_{1}\subset\Lambda_{2}$ for some integer $M\geqslant 1$. Suppose that $\Lambda_{1}$ contains $d$ linearly independent vectors in $H,H]^{d}$ for some $H>0$. Then, $\Lambda_{2}$ admits a basis where each basis vector has Euclidean norm $\leqslant d^{3/2}MH$.
@@
Let $B\subset\mathbb{R}^{d}$ be the unit ball for the Euclidean norm. Let $0<\lambda_{1}\leqslant\cdots\leqslant\lambda_{d}$} be the successive minima [^12] See [taovu, Definition 3.29] for the definition of successive minima. of $B$ with respect to $\Lambda_{2}$. Since $Mv_{1},\ldots,Mv_{d}\in\Lambda_{2}$ are linearly independent, we deduce from the above that $\lambda_{d}\leqslant\sqrt{d}MH$.
By Mahler’s theorem (see [taovu, Theorem 3.34]), we conclude that $\Lambda_{2}$ admits a basis of vectors of Euclidean norm at most $d\lambda_{d}\leqslant d^{3/2}MH$, which is what we needed to show. ∎
@@ 3.5. Short basis vectors
Then, with probability $1+O\big(d^{-d}\big)$, the lattice
$\mathcal{L}:=\Big\{(e_{1},\ldots,e_{d},f_{1},\ldots,f_{r})\in\mathbb{Z}^{d+r}\ :\ \prod_{i=1}^{d}\mathbf{b}_{i}^{e_{i}}\prod_{i=1}^{r}\mathbf{x}_{i}^{f_{i}}\equiv 1\!\!\pmod{N}\Big\}$
@@
The constant $42$ in the exponent is by no means the limit our techniques. For example, using smooth cutoffs in Definition˜3.6 would significantly reduce the error term in Lemma˜3.8 and hence lower this constant. We have not performed such optimisations to keep the paper as simple as possible.
Proof of Theorem˜3.18.
Let $M:={M_{*}}(N)$ be the integer defined in Lemma˜3.4. We introduce the auxiliary lattice
@@
By Lemmas˜3.7 and 3.8, we have the lattice point estimate
@@
Applying Proposition˜3.14, we deduce that, for all $H\geqslant e^{10d}$,
@@
Let $H_{0}:=d\lceil e^{31(d+r)}\rceil$ and $H_{1}:=\lceil d(d+r)(5/2)^{d+r}\rceil H_{0}$. We apply ˜23 and 24 twice, once with $H=H_{0}$ and once with $H=H_{1}$ to obtain the following: with probability $1+O\big(d^{-d}\big)$}, the two estimates
@@
We now apply Lemma˜3.16. By our choice of $H_{0}$ and $H_{1}$, the inequality in the statement is satisfied provided that $d$ is sufficiently large. We can assume that $d$ is large enough as, for $d\ll 1$, we have $N\ll 1$ and Theorem˜3.18 is trivially true. Thus, Lemma˜3.16 implies that $\mathcal{L}_{M}\cap[-H_{1},H_{1}]^{d+r}$} contains $d+r$ linearly independent vectors, with probability $1+O\big(d^{-d}\big)$}. By Lemma˜3.17, since $M\mathcal{L}\subset\mathcal{L}_{M}$, we conclude that, with probability $1+O\big(d^{-d}\big)$}, $\mathcal{L}$ admits a basis of vectors of Euclidean norm at most
@@
using the bound $M\leqslant e^{10d}$ given by Lemma˜3.4. This completes the proof. ∎
@@ 4. Applications to quantum computing
In this section, we prove the correctness of efficient quantum algorithms for factoring and for the discrete logarithm problem by applying our version of Regev’s number-theoretic conjecture, Theorem˜3.18.
@@ 4.1. Preparatory lemmas
This is similar to [regev, p5] or [MIT, Lemma 12] but for general values of $m$. Without loss of generality, assume that $d$ is a power of $2$, say $d=2^{l}$. We proceed to compute this product in a binary tree fashion. Let $T(k)$ be the complexity of multiplying any $k$ of these numbers $a_{i}^{t_{i}}$ modulo $N$. Note that $T(2k)\leqslant 2T(k)+O(M(mk))$ where $M(x)$ is the time needed to multiply two integers having at most $x$ bits. By the work of Harvey and van der Hoeven [annalsmult], it is known that $M(x)=O(x\log x)$, which leads to the bound
@@
This well-known fact is explained in Section A.1 of the full version of [MIT].∎
@@
Let $\mathbf{n}_{1},\ldots,\mathbf{n}_{k}$ be i.i.d. random variables uniformly distributed in $\{1,\ldots,X\}$. Then, the probability that at least $d$ of these $\mathbf{n}_{i}$ are prime numbers not dividing $N$ is $1+O(1/N)$.
@@
$\mathbb{P}\bigg(\bigcup_{\begin{subarray}{c}I\subset[k]\\ |I|>k-d\end{subarray}}\bigcap_{i\in I}E_{i}\bigg)\leqslant\sum_{l<d}\binom{k}{l}\mathopen{}\mathclose{{\left(1-\frac{c}{\log X}}}\right)^{k-l}\leqslant dk^{d}e^{-c(k-d)/\log X}\ll e^{-d^{2}}$
@@ 4.2. The discrete logarithm problem
We now prove Theorem˜1.2. For convenience, we restate it here.
@@
$\mathcal{L}_{\underline{\mathbf{b}},g,y}:=\Big\{(e_{1},\ldots,e_{d},f_{1},f_{2})\in\mathbb{Z}^{d+2}\ :\ \Bigg(\prod_{i=1}^{d}\mathbf{b}_{i}^{e_{i}}\Bigg)g^{f_{1}}y^{f_{2}}\equiv 1\!\!\pmod{N}\Big\}$.
By Theorem˜3.18, with probability $1+O(d^{-d})$, this lattice has a basis consisting of vectors of Euclidean norm $\ll e^{42d}$.
Computing the discrete logarithm of $y$ with respect to the base $g$ reduces (with a polynomial-time classical algorithm) to computing a short basis for $\mathcal{L}_{\underline{\mathbf{b}},g,y}$. To see why this is the case, suppose that we managed to compute a basis $v_{1},\ldots,v_{d+2}$ for $\mathcal{L}_{\underline{\mathbf{b}},g,y}$ with $\max_{i}\mathopen{}\mathclose{{\left\|v_{i}}}\right\|_{2}\ll e^{42d}$. It is then easy to find a vector in $\mathcal{L}_{\underline{\mathbf{b}},g,y}$ of the form $(0,\ldots,0,x,1)$ for some integer $x$ (note that such a vector exists since we assume that $y\in\mathopen{}\mathclose{{\left\langle g}}\right\rangle$). Indeed, this amounts to solving a linear system with integer coefficients. The complexity of solving an integer linear system is polynomial in the dimensions of the matrix and the number of bits of the coefficients, which are both $O(d)$ since $\max_{i}\mathopen{}\mathclose{{\left\|v_{i}}}\right\|_{2}\ll e^{42d}$ (see [bareiss]). This yields an integer $x$ such that $g^{x}\equiv y\pmod{N}$.
@@
    Generate $d$ primes $b_{1},\ldots,b_{d}$ independently and uniformly at random in the set of primes $\leqslant X$ not dividing $N$. To do this, it suffices to generate $d^{4}$ independent random integers $\leqslant X$. By Lemma˜4.3, the required conditions will be satisfied for at least $d$ of those with probability $1+O(1/N)$. This first step takes polynomial time on a classical computer using the AKS primality test [aks].
@@
    Use the procedure described by Ekerå and Gärtner [discretelogs] to obtain a basis for the lattice $\mathcal{L}_{\underline{b},g,y}$. This involves making $O(d)$ calls to a quantum circuit, followed by polynomial-time classical post-processing. This step is now guaranteed to succeed with probability $\Theta(1)$, using the fact that $\mathcal{L}_{\underline{b},g,y}$ has a basis of vectors of Euclidean norm $\ll e^{42d}$ with probability $1+O(d^{-d})$ (Theorem˜3.18).
@@
The only modification needed to the analysis of the quantum circuit in [discretelogs] comes from the fact that $b_{1},\ldots,b_{d}$ are not quite as small as in [discretelogs]. In [discretelogs] (and more generally in [MIT, regev]), the primes $b_{i}$ are assumed to have $O(\log d)$ bits. In the present situation, we instead have the bound $b_{i}\leqslant X=d^{10^{3}d}$, i.e. each $b_{i}$ has $O(d\log d)$ bits.
The only place in [discretelogs] where the assumption on the size of the $b_{i}$’s is used is in [discretelogs, Lemma 3]. In turn, the only point in the proof of [discretelogs, Lemma 3] where this assumption is needed is when [MIT, Lemma 5] is invoked (as a black box). The proof of [MIT, Lemma 5] is given in [MIT, Section 5], and the size assumption on the $b_{i}$’s only comes up in [MIT, Lemma 12].
The result [MIT, Lemma 12] essentially [^14] There are extra details pertaining to the precise use of qubits (e.g. restoring the ancilla qubits to $|0\rangle$), but these will be the same in our setup. states that there is a quantum circuit using $O(d\log^{3}d)$ gates and $O(d\log^{3}d)$ qubits to perform the computation of $\prod_{i=1}^{d}a_{i}^{t_{i}}$ where $t_{i}\in\{0,1\}$ and $a_{i}$ are integers on $O(\log d)$ bits. This is a special case of Lemma˜4.1 (used together with Lemma˜4.2) applied with $m=O(\log d)$. In our case, we just apply Lemma˜4.1 with $m=O(d\log d)$, together with Lemma˜4.2 to convert the classical circuit into a quantum one. This yields a quantum circuit having $O(d^{2}\log^{3}d)$ gates and $O(d^{2}\log^{3}d)$ qubits to compute $\prod_{i=1}^{d}b_{i}^{t_{i}}$.
The number of gates of the quantum circuit [MIT, Lemma 5] is
@@
(because [MIT, Lemma 12] is used $O(d)$ times, see [MIT, Algorithm 5.2] and the surrounding explanations). Note that this is the same as for Regev’s algorithm (see [regev, p5]). In our case, the number of gates is
@@
The space optimisations of Ragavan and Vaikuntanathan keep the total number of qubits for their circuit [MIT, Lemma 5] under
@@
due to the way the qubits are used and restored in the main loop of [MIT, Algorithm 5.2]. In our situation, the number of qubits is
@@
In summary, our final quantum circuit has $O(n^{3/2}\log^{3}n)$ gates and $O(n\log^{3}n)$ qubits. As noted in [discretelogs], the fact that the two elements $g$ and $y$ are not small ($g$ and $y$ can be as large as $N$, as opposed to the $b_{i}$’s) does not affect these complexity bounds. ∎
@@
In the proof of Theorem˜1.2, we used the naive Lemma˜4.2 to convert a classical circuit into a quantum one. As mentioned in Remark˜1.3, it is possible to save a factor of $\log n$ in the space cost for Theorems˜1.2 and 1.1 by reusing certain qubits when performing the quantum computation corresponding to Lemma˜4.1.